# On uniqueness of radial potentials for given Dirichlet spectra with distinct angular momenta

### v1

@@ 7 Uniqueness result in the case $\ell=0$ and $\ell=1$
In particular, this implies that the spectral map associated with the pair $(\ell_{1},\ell_{2})=(0,1)$ is injective.
8 Uniqueness result in the case $\ell=0$ and $\ell=2$
In this section, we assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Proposition 5.1 for both $\ell=0$ and $\ell=2$. As before, the condition for $\ell=0$ implies that $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$. For $\ell=2$, we set $g:=T_{2}[\zeta]$, and invoke the corresponding symmetry result from Theorem 6.6, which tells us that
$A_{2}(D)\bigl[g(x)\bigr]$
is odd in the sense of distributions, where $D=\frac{\partial}{\partial x}$ and $A_{2}(D)$ is the differential operator associated to the polynomial
$A_{2}(t)=\frac{t^{2}}{4}-\frac{3t}{2}+3.$
Our goal is again to deduce from this a differential equation satisfied by $\zeta$, and eventually to conclude that $\zeta=0$ under these assumptions. The method follows the same strategy as in the previous section: we express the symmetry condition as an ODE involving $\zeta$, analyze its parity structure, and reduce the problem to a uniqueness statement for solutions of a singular differential equation on $(0,1)$.
Let us set $f:=S_{1}[\zeta]$, so that $g=-S_{2}[f]$. We now differentiate the expression $A_{2}(D)\bigl[g(x)\bigr]$ four times. Since the application of an even number of derivatives preserves parity in the distributional sense, and since the operator $A_{2}(D)$ commutes with differentiation, it follows that
@@ 8 Uniqueness result in the case $\ell=0$ and $\ell=2$
is odd, where we have used Lemma 4.1 (vi) with $\ell=2$.
Using again Lemma 4.1 (vi), now with $\ell=1$, and substituting into the above expression, we obtain that the function $A_{2}(D)\bigl[F(x)\bigr]$ is odd in the sense of distributions, where
@@
As in the cases $\ell=0$ and $\ell=1$, we now define $y(x):=\zeta^{\prime}(x)$. Since $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$, it follows that $y$ is an even function.
Applying the differential operator $4A_{2}(D)=D^{2}-6D+12$ to the function $F(x)$, we define
$G(x):=4A_{2}(D)[F(x)]=\left(\frac{d^{2}}{dx^{2}}-6\frac{d}{dx}+12\right)\left(y^{(3)}+\frac{12}{x}\,y^{\prime\prime}+\frac{24}{x^{2}}y^{\prime}-\frac{24}{x^{3}}y\right).$
We find:
@@
$+\left(-6+\frac{12}{x}\right)y^{(4)}(x)+y^{(5)}(x).$ (8.1)
@@
More precisely, we obtain the following lemma :
Lemma 8.1.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Lemma 5.1 for $\ell=0$ and $\ell=2$, and set $y(x):=\zeta^{\prime}(x)$. Then $y$ satisfies the differential equation
$\left(\frac{1}{x}+\frac{1}{1-x}-1\right)y^{(4)}(x)-6\left(\frac{1}{x}-\frac{1}{1-x}\right)y^{(3)}(x)$
$\left[12\left(\frac{1}{x}+\frac{1}{1-x}\right)-8\left(\frac{1}{x^{3}}+\frac{1}{(1-x)^{3}}\right)-6\left(\frac{1}{x^{2}}+\frac{1}{(1-x)^{2}}\right)\right]y^{\prime\prime}(x)$
$\left[24\left(\frac{1}{x^{2}}-\frac{1}{(1-x)^{2}}\right)+24\left(\frac{1}{x^{4}}-\frac{1}{(1-x)^{4}}\right)+36\left(\frac{1}{x^{3}}-\frac{1}{(1-x)^{3}}\right)\right]y^{\prime}(x)$
$\left[24\left(\frac{1}{x^{3}}+\frac{1}{(1-x)^{3}}\right)+24\left(\frac{1}{x^{5}}+\frac{1}{(1-x)^{5}}\right)+36\left(\frac{1}{x^{4}}+\frac{1}{(1-x)^{4}}\right)\right]y(x)=0$
in the sense of distributions on $(0,1)$.
The indicial equation of this differential equation has four solutions, namely $\rho=-2,\,1,\,3,\,4$. As before, we must discard the solution whose asymptotic behavior near $x=0$ is of the form $x^{-2}$.
We shall prove below that there exists a unique even solution in $L^{2}(0,1)$ to this equation. The argument proceeds as follows. Assume, by contradiction, that there exist two linearly independent even solutions. Then, by taking a suitable linear combination, we could find an even solution $y(x)$ such that
$y(0)=y^{\prime}(0)=y^{\prime\prime}(0)=0.$
We now apply the same idea as before, making use of the relation
[…]
which we recompute by performing several integrations by parts. By parity, we first observe that
$\int_{0}^{1}y^{(5)}(x)\,y(x)\,dx=0.$
Moreover, we have the following identities:
$\int_{0}^{1}\left(-6+\frac{12}{x}\right)y^{(4)}(x)\,y(x)\,dx=\int_{0}^{1}\left(-6+\frac{12}{x}\right)y^{\prime\prime}(x)^{2}\,dx\;-\;48\int_{0}^{1}x^{-3}\,y^{\prime}(x)^{2}\,dx\;+\;144\int_{0}^{1}x^{-5}\,y(x)^{2}\,dx,$
$\int_{0}^{1}\left(12-\frac{72}{x}\right)y^{(3)}(x)\,y(x)\,dx=108\int_{0}^{1}x^{-2}\,y^{\prime}(x)^{2}\,dx\;-\;216\int_{0}^{1}x^{-4}\,y(x)^{2}\,dx,$
and
$\int_{0}^{1}P(x)\,y^{\prime\prime}(x)\,y(x)\,dx=-\int_{0}^{1}P(x)\,y^{\prime}(x)^{2}\,dx\;+\;\frac{1}{2}\int_{0}^{1}P^{\prime\prime}(x)\,y(x)^{2}\,dx,$
where
$P(x)\;:=\;-\frac{96}{x^{3}}-\frac{72}{x^{2}}+\frac{144}{x},\qquad P^{\prime\prime}(x)\;=\;-\frac{1152}{x^{5}}-\frac{432}{x^{4}}+\frac{288}{x^{3}}.$
Furthermore,
$\int_{0}^{1}\left(\frac{288}{x^{4}}+\frac{432}{x^{3}}+\frac{288}{x^{2}}\right)y^{\prime}(x)\,y(x)\,dx=\int_{0}^{1}\left(\frac{576}{x^{5}}+\frac{648}{x^{4}}+\frac{288}{x^{3}}\right)y(x)^{2}\,dx.$
Combining all the above identities, we obtain the following compact expression after simplification:
[…]
Now, we make use of the weighted Hardy inequalities in the form
$\int_{0}^{1}\frac{y(x)^{2}}{x^{5}}\,dx\;\leq\;\frac{1}{4}\int_{0}^{1}\frac{y^{\prime}(x)^{2}}{x^{3}}\,dx\,,\qquad\int_{0}^{1}\frac{y(x)^{2}}{x^{4}}\,dx\;\leq\;\frac{4}{9}\int_{0}^{1}\frac{y^{\prime}(x)^{2}}{x^{2}}\,dx\,.$
As a consequence, we deduce that
[…]
We start from (8.2) and observe as above that, for $x\in(0,\frac{1}{2}]$, we have
$f(x)\;:=\;\frac{12}{x^{3}}+\frac{84}{x^{2}}-\frac{144}{x}\geq 96+336-288=144\,.$
Therefore,
$\int_{0}^{1}f(x)\,y^{\prime}(x)^{2}\,dx\;\geq\;144\int_{0}^{\frac{1}{2}}y^{\prime}(x)^{2}dx+\int_{\frac{1}{2}}^{1}f(x)y^{\prime}(x)^{2}dx=\int_{\frac{1}{2}}^{1}(f(x)+144)\,y^{\prime}(x)^{2}dx\,$,
where we have used the symmetry of $y^{\prime}(x)^{2}$ for the last equality. It remains to verify that $f(x)+144$ is positive on $(0,1]$ but its infimum is at $x=1$ and equals $96$.
Since all the integral in (8.2) are positive, we conclude that $y$ must be identically zero. In particular, we have just shown that there exists only one even solution $y(x)\in L^{2}(0,1)$.
The use of a computer is required to determine the unique odd solution $\zeta\in L^{2}(0,1)$ described in Lemma 8.1, (recall that $y(x)=\zeta^{\prime}(x)$). It is explicitly given by
$\zeta(x)\;=\;A\left(4x^{3}-6x^{2}-2x+2-\frac{3(2x-1)}{2\,(x^{2}-x+1)^{2}}+\frac{2(2x-1)}{x^{2}-x+1}\right),$
where $A\in\mathbb{R}$ is an arbitrary constant. Note that the expression $x^{2}-x+1$ is invariant under the transformation $x\mapsto 1-x$.
Unfortunately, this is only a necessary condition, because we applied the operator $A_{2}(D)$ to $T_{2}(\zeta)$ after differentiating four times. We must still check, for this specific $\zeta$, whether $A_{2}(D)\bigl[T_{2}(\zeta)\bigr]$ is indeed an odd function in the sense of distributions.
Recall that, following Rundell and Sacks [23, after Lemma 3.2], the operator $T_{2}$ applied to a function $\zeta$ is defined by
$T_{2}(\zeta)(x)=-\,\zeta(x)-12\,x\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt+24\,x^{3}\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt\,.$
A direct computation yields
$4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr](x)=(288\,x^{3}-432\,x^{2}+144\,x)\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt+(-144\,x+72)\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt$
$\zeta^{\prime\prime}(x)+6\,\zeta^{\prime}(x)-12\,\zeta(x)-\frac{12}{x}\,\zeta^{\prime}(x)+\frac{72}{x}\,\zeta(x)-\frac{48}{x^{2}}\,\zeta(x)\,.$
Evaluating at $x=\tfrac{1}{2}$ (with $A=1$), all integral terms vanish and we obtain, since $\zeta$ is odd,
@@
$4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr]\Big|_{x=\frac{1}{2}}=-60\,\zeta\!\left(\tfrac{1}{2}\right)-18\,\zeta^{\prime}\!\left(\tfrac{1}{2}\right)-\zeta^{\prime\prime}\!\left(\tfrac{1}{2}\right)=-18\,\zeta^{\prime}\!\left(\tfrac{1}{2}\right)=90\,.$
Since $4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr](x)$ is a smooth function, if it were odd with respect to the midpoint $x=\tfrac{1}{2}$, its value at $x=\tfrac{1}{2}$ would necessarily vanish. Therefore, the only possible choice ensuring the oddness of $4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr]$ in the sense of distributions is to take $A=0$.
We thus obtain a completeness result in the case $(\ell_{1},\ell_{2})=(0,2)$, in full analogy with the case $(\ell_{1},\ell_{2})=(0,1)$:

### v3

@@ 7 Uniqueness result in the case $(\ell_{1},\ell_{2})=(0,1)$
In particular, this implies that the differential spectral map associated with the pair $(\ell_{1},\ell_{2})=(0,1)$ is injective.
8 Uniqueness result in the case $(\ell_{1},\ell_{2})=(0,2)$
Throughout this subsection, we assume that $\zeta$ satisfies the hypotheses of Proposition 5.1 for both $\ell=0$ and $\ell=2$ and we aim to show that $\zeta$ vanishes almost everywhere on $(0,1)$ . As before, the condition for $\ell=0$ implies that $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$ and for $\ell=2$ , it means by Theorem 6.6 that the function $A_{2}(D)\bigl[T_{2}[\zeta]\bigr]$ is odd in the sense of distributions.
In order to prove that $\zeta$ vanishes almost everywhere on $(0,1)$, we follow the same strategy as in the previous Subsection. The aim is to prove that $y=\zeta^{\prime}$ is a solution of a linear fourth-order differential equation on $(0,1)$ and satisfies the conditions
$y\!\left(\tfrac{1}{2}\right)=y^{\prime}\!\left(\tfrac{1}{2}\right)=y^{\prime\prime}\!\left(\tfrac{1}{2}\right)=y^{(3)}\!\left(\tfrac{1}{2}\right)=0\,.$
It then follows that $y$ vanishes identically on $(0,1)$ , and consequently $\zeta$ vanishes almost everywhere on $(0,1)$ , since it is an odd function.
Let us set $f:=S_{1}[\zeta]$ and $g=-S_{2}[f]$, so that $g=T_{2}[\zeta]$ . We now differentiate the expression $A_{2}(D)\bigl[g\bigr]$ four times. Since the application of an even number of derivatives preserves parity in the distributional sense, and since the operator $A_{2}(D)$ commutes with differentiation, it follows that
@@ 8 Uniqueness result in the case $(\ell_{1},\ell_{2})=(0,2)$
is odd, where we have used Lemma 4.1 (iv) with $\ell=2$ . Using again Lemma 4.1 (iv), now with $\ell=1$ , and substituting into the above expression, we obtain that the function $A_{2}(D)\bigl[F\bigr]$ is odd in the sense of distributions, where for all $x\in(0,1)$ ,
@@
As in the cases $\ell=0$ and $\ell=1$, we now define $y(x):=\zeta^{\prime}(x)$ . Since $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$, it follows that $y$ is an even function. Applying the differential operator $4A_{2}(D)=D^{2}-6D+12$ to the function $F$, we define
$G(x):=4A_{2}(D)[F(x)]=\left(\frac{d^{2}}{dx^{2}}-6\frac{d}{dx}+12\right)\left(y^{(3)}+\frac{12}{x}\,y^{\prime\prime}+\frac{24}{x^{2}}y^{\prime}-\frac{24}{x^{3}}y\right)$
and we find:
@@
$+\left(-6+\frac{12}{x}\right)y^{(4)}(x)+y^{(5)}(x)\,.$
@@
We now prove that $y\!\left(\tfrac{1}{2}\right)=y^{\prime}\!\left(\tfrac{1}{2}\right)=y^{\prime\prime}\!\left(\tfrac{1}{2}\right)=y^{(3)}\!\left(\tfrac{1}{2}\right)=0$ . First, since $y$ is even, $y^{\prime}$ and $y^{(3)}$ are both odd and $y^{\prime}\!\left(\tfrac{1}{2}\right)=y^{(3)}\!\left(\tfrac{1}{2}\right)=0$ . Recall that (see (3.9) in [27]), for all $x\in(0,1)$ ,
$g(x)=T_{2}[\zeta](x)=-\zeta(x)-12x\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt+24x^{3}\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt\,.$
Applying the differential operator $4A_{2}(D)=D^{2}-6D+12$ to $g$, we obtain for all $x\in(0,1)$ ,
$4A_{2}(D)[g](x)=$ $-\zeta^{\prime\prime}(x)+\left(6-\frac{12}{x}\right)\zeta^{\prime}(x)+\left(-\frac{48}{x^{2}}+\frac{72}{x}-12\right)\zeta(x)$
$+72(1-2x)\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt+144x(x-1)(2x-1)\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt\,.$
Evaluating this expression at $x=\tfrac{1}{2}$ and using that $\zeta$ is odd and $y=\zeta^{\prime}$, we obtain
$4A_{2}(D)\bigl[g\bigr]\!\left(\tfrac{1}{2}\right)=-\zeta^{\prime\prime}\!\left(\tfrac{1}{2}\right)-18\zeta^{\prime}\!\left(\tfrac{1}{2}\right)-60\zeta\!\left(\tfrac{1}{2}\right)=-18\,y\!\left(\tfrac{1}{2}\right)\,.$
@@
Since $4A_{2}(D)\bigl[g\bigr](x)$ is odd, we therefore conclude that $y\!\left(\tfrac{1}{2}\right)=0$ .
Proceeding in the same way, we compute
$4D^{2}A_{2}(D)[g](x)=$ $-\zeta^{(4)}(x)+\left(6-\frac{12}{x}\right)\zeta^{(3)}(x)+\left(-\frac{24}{x^{2}}+\frac{72}{x}-12\right)\zeta^{\prime\prime}(x)$
$+\left(\frac{24}{x^{3}}+\frac{216}{x^{2}}-\frac{144}{x}\right)\zeta^{\prime}(x)+\left(\frac{288}{x^{3}}-\frac{576}{x^{2}}\right)\zeta(x)$
$+864(2x-1)\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt\,.$
Recalling that $\zeta$ is odd, $y=\zeta^{\prime}$ and $y\!\left(\tfrac{1}{2}\right)=0$ , we evaluate at $x=\tfrac{1}{2}$ and obtain
$4\,D^{2}A_{2}(D)\bigl[g\bigr]\!\left(\tfrac{1}{2}\right)=-\,y^{(3)}\!\left(\tfrac{1}{2}\right)-18\,y^{\prime\prime}\!\left(\tfrac{1}{2}\right)+36\,y^{\prime}\!\left(\tfrac{1}{2}\right)+768\,y\!\left(\tfrac{1}{2}\right)=-18\,y^{\prime\prime}\!\left(\tfrac{1}{2}\right)\,.$
Since $4\,D^{2}A_{2}(D)\bigl[g\bigr]$ is also an odd function, we conclude, as above, that $y^{\prime\prime}\!\left(\tfrac{1}{2}\right)=0$ .
In conclusion, $y$ satisfies a fourth–order linear differential equation with the initial conditions
$y\!\left(\tfrac{1}{2}\right)=0,\qquad y^{\prime}\!\left(\tfrac{1}{2}\right)=0,\qquad y^{\prime\prime}\!\left(\tfrac{1}{2}\right)=0,\qquad y^{(3)}\!\left(\tfrac{1}{2}\right)=0\,.$
The Cauchy–Lipschitz theorem then implies that $y\equiv 0$. Since $y=\zeta^{\prime}$ and $\zeta$ is odd, this in turn forces $\zeta\equiv 0$ . We thus obtain a completeness result in the case $(\ell_{1},\ell_{2})=(0,2)$, in full analogy with the case $(\ell_{1},\ell_{2})=(0,1)$: